rvc (rvc):

A clock regulated by a seconds pendulum keeps correct time.During summer the length of the pendulum increases to 1.01m.How much will the clock gain or lose in one day?

rvc (rvc):

@Mashy @nirmalnema @praxer

rvc (rvc):

@cwrw238 help

rvc (rvc):

@paki

rvc (rvc):

can u help

OpenStudy (anonymous):

can u tell whether the clock will lose or gain?

rvc (rvc):

no idea

rvc (rvc):

@ganeshie8

OpenStudy (anonymous):

@amistre64 can you do some help here??

OpenStudy (amistre64):

im thinking that a pendulum swing regulates itself and that the time frame essentially remains constant.

OpenStudy (amistre64):

but im no expert :)

rvc (rvc):

@Mashy help pls

OpenStudy (anonymous):

The period of a simple pendulum is proportional to the square root of its length. If you assume the period depends on the length and the acceleration due to gravity, then the result follows from dimensional analysis. Knowing that period is proportional to square root of length, you should be able to work out the change in period for a 1percent change in length.

rvc (rvc):

@Mashy help yaar

rvc (rvc):

@ProfBrainstorm help

OpenStudy (anonymous):

yes.. @ProfBrainstorm helped u enough.. the length went from 1 to 1.01 so the change = 1.01 - 1 = 0.01 so the percent increase = 0.01 * 100 = 1 % also you know \[T = 2 \pi \sqrt{ \frac{L}{g}}\]

OpenStudy (anonymous):

so how much is the percentage increase in Time ?!

rvc (rvc):

hey how u know L=1m

OpenStudy (anonymous):

seconds pendulum = L = 1m

OpenStudy (anonymous):

Just assume that the formula for the period of a simple pendulum is valid and plug in values for L and g to find T. L=1.01m and g=9.81m/s^2 Then you can work out how many oscillations will have occured in one day. Note that the period of an accurate seconds pendulum is actually 2 seconds, so it marks one second on each swing.